### v2

# stanhf: HistFactory models in the probabilistic programming language Stan

Abstract
In collider physics, experiments are often based on counting the numbers of events in bins of a histogram. We present a new way to build and analyze statistical models that describe these experiments, based on the probabilistic programming language Stan and the HistFactory specification. A command-line tool transpiles HistFactory models into Stan code and data files. Because Stan is an imperative language, it enables richer and more detailed modeling, as modeling choices defined in the HistFactory declarative specification can be tweaked and adapted in the Stan language. Stan was constructed with automatic differentiation, allowing modern computational algorithms for sampling and optimization.
I Introduction
Interpreting the results of a physics experiment requires a statistical model. This model describes the relationship between a theory and potential observations, including statistical and systematic errors that could affect the outcome. In collider physics, experiments are often based on counting the numbers of events in bins of a histogram. We present a new way to build and analyze these statistical models, based on the probabilistic programming language (PPL) Stan [1, 2]. These computer representations of a statistical model preserve and communicate an experimental analysis and enable it to be modified or recasted.
More generally, PPLs are languages that are designed for statistical modeling (see e.g., refs. [3, 4]). Like most languages, they support random number generation and their outputs can be probabilistic; significantly, however, they are designed to make it easy to express statistical models and perform statistical inference. Besides Stan, examples of PPLs include WinBUGS [5], Turing.jl [6], PyMC [7], Edward [8, 9], pyro/numpyro [10, 11], WinBUGS [5] and RooFit/RooStats [12, 13, 14]. In the context of high-energy physics, PPLs were recently explored in refs. [15, 16, 17, 18, 19].
To understand the form of statistical models for these counting experiments, we start from the Poisson likelihood for data binned in a histogram or several histograms
where the index runs over all histogram bins, and and represent the observed and expected numbers of events in bin , respectively. There are several contributions to the expected number of events in a bin from specific backgrounds and from potential signals, such that we write
The nominal rates for each contribution to each bin, , can be affected by systematic errors. These errors are modeled by introducing nuisance parameters . These parameters control multiplicative factors and additive terms that modify the nominal rates,
where the factor represents the product of all multiplicative systematic effects and the term represents the sum of all additive systematic effects for the nominal rate for bin contribution . Thus, they can be broken down into individual systematic effects,
where indexes individual systematic effects, and and represent individual multiplicative and systematic effects for bin , contribution and effect . Lastly, the parameters might be constrained by prior knowledge or auxiliary experiments. Thus, combining these changes to Eq. 1, we arrive at
where represents prior knowledge or data from auxiliary measurements.
On the other hand, rather than binning events into histograms, one could model the events using a probability density function (pdf). Suppose that we observed events with a discriminating variable, e.g. invariant mass, . Rather than binning with respect to the discriminating variable as in Eq. 1, we write
where the pdf models the discriminating variable for each event. Each observed event is marked by its value of the discriminating variable, and thus Eq. 6 is known as a marked Poisson model.
Previously, the HistFactory specification was developed to represent binned models [20]. This is a declarative specification of a statistical model of the form Eq. 5 — the modifiers and auxiliary term in Eq. 5 can be built from a set of predeclared choices. The specification was written in xml language and was intended to be interpreted by the RooFit/RooStats framework based on ROOT histogram functionality [12, 13, 14]. More recently, a json schema was designed to express HistFactory models in json [21, 22].
At the same time, the Python package pyhf [23, 24, 25] was developed to interpret and analyze HistFactory models from json files. As Python is somewhat ubiquitous compared to ROOT, this increased the portability and potentially increases the longevity of these statistical models. pyhf, furthermore, takes advantage of automatic differentiation, so that derivatives are available for optimization. The results of searches for new particles at the Large Hadron Collider (LHC) are often preserved as HistFactory models and hosted publicly on the HEPData webpage [26]. At present, HistFactory models and pyhf are used by the ATLAS, MicroBooNE [27], Belle and Belle-II [28], and MODE [29] collaborations. Bayesian analyses are possible through an interface to PyMC [7] in bayesian_pyhf [30].
We present a complementary tool to the successful pyhf and HistFactory based on Stan [1, 2]. Stan is an imperative probabilistic programming language with automatic differentiation [31], enabling models to be analyzed in a Bayesian framework using Hamiltonian Monte Carlo (HMC; see e.g. refs. [32, 33]). This is an efficient algorithm for obtaining a sample-representation of the posterior distribution. Stan started development in 2011, and is used in industry [34, 35] and academia, including cognitive, social and political sciences, medical statistics, and astrophysics [36, 37, 38]. As we build on HistFactory, we focus on binned models, though unbinned models such as Eq. 6 could be programmed in Stan.
Because Stan is an imperative language, it is more flexible and expressive than declarative specifications. This enables richer and more detailed modeling, as the modeling choices defined in the HistFactory declarative specification can be tweaked and adapted in the Stan language. This is a major difference from pyhf and bayesian_pyhf. Although Stan was designed primarily for Bayesian modeling and computation, frequentist statistics, such as confidence intervals and p-values, are available in stanhf through an interface with pyhf. Stan aims to offer state-of-the-art performance and supports computation on GPUs [39]. Performance comparisons to JAX [40], a possible computational backend in pyhf, indicate that Stan outperforms JAX on the CPU but that JAX outperforms Stan on the GPU [41].
II Quickstart
The package and dependencies can be installed by
The source code can be obtained by
You can obtain a HistFactory model in json specification from HEPData. E.g., using curl,
or just browsing the landing page for the resource e.g. here. This downloads and extracts files, including
for HistFactory models for searches for squarks by ATLAS at the LHC in ref. [42].
The HistFactory model RegionA/BkgOnly.json can be transpiled into Stan using the command line interface (CLI),
This writes a HistFactory model as a Stan program, compiles it, and validates it. If necessary, cmdstan [43] is automatically downloaded and installed so that Stan models can be compiled. These steps, as well as metadata about the model, are written to stdout:
As indicated, three files are written to disk, and the Stan model is compiled into an executable:
The process is summarized for simplified example HistFactory models in Figs. 1 and 2. The simplified models include a normalization factor (denoted by normfactor in the HistFactory specification) and a normalization systematic uncertainty (denoted by normsys in the HistFactory specification), respectively. They can be found in the stanhf source code at examples/normfactor.json and examples/normsys.json, respectively.
As suggested in the output, you can run this model by e.g.
For further information on running an analysis, see Sections IV and V.
III Further details
III.1 Structure of Stan model
Stan is a strongly typed and imperative probabilistic programming language. Stan programs, including those created by stanhf, define a statistical model through special blocks. We present example Stan programs produced from the simplified examples examples/normfactor.json and examples/normsys.json in Figs. 1 and 2, respectively.
First, the functions block. In stanhf we use this block to define functions for interpolation and constraint terms that are required to implement the HistFactory specification.
We declare the data in the data block; here, data includes experimental data, as well as nominal event numbers, and hyperparameters that define modifiers and constraint terms. We do not define data here; it is written to a separate json file. In the context of high-energy physics, the natural separation between a model and observed data in Stan makes data-blinding straight-forward [44], as observed data can be substituted without altering the model or model file.
Occasionally, we need to make transformations of the data before we begin sampling, e.g., to aggregate uncertainties for a constraint term.
Unknown parameters, associated with modifiers, are declared in the parameters block. These parameters are marginalized or profiled in an analysis.
The expected event rates are computed by applying modifiers to the nominal rates in the transformed parameters block.
The Poisson likelihood and constraint terms for background knowledge or auxiliary measurements are entered in the model block.
Finally, we simulate data from the model in the generated quantities block. This simulated data can be used to compute posterior predictive distributions.
III.2 Choices of prior
Stan is agnostic about whether auxiliary terms in Eq. 5 are part of the prior or likelihood, since it is only their product that matters. Thus, we can interpret stanhf models as having implicit flat priors for parameters between any bounds placed on them in the HistFactory model, and constraint terms that are part of the likelihood.
If we desired other choices of prior, we could code them in the Stan language and rebuild the model. E.g., for a Gaussian prior for the parameter mu, , we could add
to the model block. These choices cannot be automated and should be considered by a user.
Lastly, we may wish to choose a conjugate prior [45] — that is, a prior such that the prior and posterior belong to the same family of distributions. In this case, we may update using the auxiliary data as a preliminary step and specify only the posterior in the Stan model. That is, we could remove the constraint term and replace it with a posterior .
III.3 Command-line options
The behavior of the stanhf program can be altered through command-line options. E.g., the compilation and validation steps may be disabled. To see these flags, use the --help flag:
III.4 Patches
Patches are a way to modify a HistFactory model by e.g. adding a specific signal. They are supported in stanhf by an optional argument patch. E.g., we can patch the model RegionA/Bkgonly.json from Section II
This applies the first patch in the RegionA/patchset.json collection of patches to RegionA/Bkgonly.json and converts the resulting model to Stan. We use pyhf to apply patches.
III.5 Null parameters and modifiers
To simplify models, stanhf ignores modifiers and parameters that cannot influence the expected numbers of events. This occurs when, e.g., a modifier varies between lo and hi, but lo == hi. These modifiers are reported as being null. Parameters that are associated only with null modifiers are themselves null.
III.6 Validation and testing
The stanhf interpretation of a HistFactory model is validated against pyhf by
1.
Checking that the parameter names and sizes agree; see stanhf.Converter.validate_par_names.
2.
Checking that the log-likelihoods agree up to a constant term. This is performed by checking that
The log-likelihoods are evaluated at randomly chosen points found by perturbing the suggested initial choices by noise. See stanhf.Converter.validate_target.
A constant term in the log-likelihood cannot impact statistical inferences currently available in pyhf or Stan and, for performance, is the default behavior in Stan.111Though this constant can be relevant in model selection if it differs between models.
As part of the stanhf testing set, these checks are performed for several models and patches available from HEPData; see e.g. stanhf/tests/test_hepdata.py. These tests, and others, can be ran using
from within the stanhf source code.
IV Bayesian inference
The stanhf CLI builds an executable Stan program and writes associated data files. We briefly describe how to use them for Bayesian inference. For concreteness, we look here at the examples/normfactor.json model shown in Fig. 1. In this model, a signal is characterized by a strength parameter and corresponds to the background-only model. First, we must build the model and sample from it,
This uses the HMC sampling algorithm with 4 chains, in conjunction with the stanhf data files. This is the cmdstan interface to Stan, where we run Stan at the command-line; you can alternatively explore language-specific interfaces. There are Python, Julia, R and MATLAB interfaces [46], as well as a web interface [47]. For all CLI options, see examples/model --help.
From the CLI interface, the output from Stan shows,
[…]
Here we demonstrate ArviZ, since it is a Python package, and most familiar to users in high-energy physics. After installing ArviZ,
you can analyze the Stan results. We can find the mean, standard deviation and credible region of the strength parameter,
We may check the effective sample size and diagnostic [53],
We can plot posterior distributions. For example, to show the one-dimensional posterior pdf
[…]
Lastly, the Bayes factor [54] and Bayes factor surface [55] can be computed using a Savage-Dickey ratio [56]. The Bayes factor surface as a function of relative to the background-only model could be found by,
[…]
This calculation assumed that we sampled the posterior using a flat prior for , though the result is independent of that choice.222Other priors are supported using the prior keyword argument. To further explore Bayesian workflows, this model can be compiled, run and analyzed online here using the Stan online playground [47].
V Frequentist inference
As well as using stanhf for Bayesian inference using the powerful HMC functionality of Stan, we can harness the expressive power of Stan models in a frequentist setting. stanhf provides classes and functions to use a Stan model and optimizer in place of a pyhf model or native pyhf optimizer in the statistical inference package of pyhf.infer, which uses results in ref. [57].333Frequentist inference is possible in Stan itself using simulation; see e.g. ref. [58].
As an example, we again use examples/normfactor.json, which declares that mu is the parameter of interest (POI). Using pyhf, we can plot the statistic [59] as a function of :
[…]
Here, mock_pyhf_backend provides a context in which pyhf uses the Stan optimizer as the backend, and MockPyhfModel creates a model that can be used in conjunction with the Stan optimizer. In this example, we see an upper limit of at for the parameter mu.
Thus, we can tweak models in Stan — going beyond the restrictions of the declarative HistFactory specification — and still access pyhf statistical machinery. In this case, however, it is a user’s responsibility to ensure that any assumptions or approximations required by asymptotic formulae in pyhf.infer are valid.
The frequentist analysis in pyhf.infer requires one to select a particular POI. To perform tests on particular values of the POI, it must be possible to fix the POI. Thus, if a POI is specified in a HistFactory model, stanhf writes Stan code that allows it to be fixed. For a POI named {poi}, this is implemented using the pattern:
VI Tweaking a model
Although the HistFactory declarative specification describes many common models, we may wish to go beyond it. For example, suppose we wish to model contributions to a signal, and , that depend on different powers of a modifier, e.g.,
This form could occur because of interference effects or because the contributions occur at different orders in perturbation theory.
We achieve this as follows. We begin from a model in which two signal contributions share a correlated normalization factor. This model can be found in the stanhf source code in examples/tweak.json. We generate a Stan model using stanhf examples/tweak.json. Because Stan is an imperative language, we can tweak the Stan code generated by stanhf as we wish. Thus, we edit the resulting tweak.stan model to match the required form Eq. 7. Specifically, we change the transformed parameters block such that the second contribution to the signal depends on the square of the normalization factor,
We show this process in full in Fig. 3.
VII Conclusions
We introduced stanhf — a new way to analyze and develop statistical models based on histogrammed data. The code transpiles HistFactory models into Stan, a probabilistic programming language. This allows state-of-the-art, scalable optimization and sampling using automatic differentiation and HMC. Because Stan is an imperative language, HistFactory modeling choices can be tweaked and adapted, as desired.
Stan is part of an ecosystem of Bayesian analysis software, allowing MCMC chains to be scrutinized for convergence and plotted, and even allowing models to be reran and analyzed online. Lastly, to ensure that the frequentist toolbox remains available, we provide an interface between Stan models and the inference machinery of pyhf.
Acknowledgements.
AF was supported by RDF-22-02-079.References
- [1] B. Carpenter, A. Gelman, M.D. Hoffman, D. Lee, B. Goodrich, M. Betancourt et al., Stan: A Probabilistic Programming Language, J. Stat. Softw. 76 (2017) .
- [2] Stan Development Team, Stan Modeling Language User’s Guide and Reference Manual, URL.
- [3] O.A. Martin, R. Kumar and J. Lao, Bayesian Modeling and Computation in Python, Chapman and Hall/CRC (2021), DOI.
- [4] J. Brehmer, Simulation-based inference in particle physics, Nat. Rev. Phys. 3 (2021) 305 [2010.06439].
- [5] D.J. Lunn, A. Thomas, N. Best and D. Spiegelhalter, WinBUGS — A Bayesian modelling framework: Concepts, structure, and extensibility, Stat. Comput. 10 (2000) 325–337.
- [6] T.E. Fjelde, K. Xu, D. Widmann, M. Tarek, C. Pfiffer, M. Trapp et al., Turing.jl: a general-purpose probabilistic programming language, ACM Trans. Probab. Mach. Learn. (2025) .
- [7] O. Abril-Pla, V. Andreani, C. Carroll, L. Dong, C.J. Fonnesbeck, M. Kochurov et al., PyMC: a modern, and comprehensive probabilistic programming framework in Python, PeerJ Comput. Sci. 9 (2023) .
- [8] D. Tran, A. Kucukelbir, A.B. Dieng, M. Rudolph, D. Liang and D.M. Blei, Edward: A library for probabilistic modeling, inference, and criticism, [1610.09787].
- [9] D. Tran, M.D. Hoffman, D. Moore, C. Suter, S. Vasudevan, A. Radul et al., Simple, Distributed, and Accelerated Probabilistic Programming in NeurIPS, 2018, URL.
- [10] E. Bingham, J.P. Chen, M. Jankowiak, F. Obermeyer, N. Pradhan, T. Karaletsos et al., Pyro: Deep Universal Probabilistic Programming, J. Mach. Learn. Res. 20 (2018) 1 [1810.09538].
- [11] D. Phan, N. Pradhan and M. Jankowiak, Composable Effects for Flexible and Accelerated Probabilistic Programming in NumPyro, [1912.11554].
- [12] I. Antcheva et al., ROOT: A C++ framework for petabyte data storage, statistical analysis and visualization, Comput. Phys. Commun. 180 (2009) 2499 [1508.07749].
- [13] W. Verkerke and D.P. Kirkby, The RooFit toolkit for data modeling in The 13th International Conference on Computing in High-Enery and Nuclear Physics (CHEP 2003), L. Lyons and M. Karagoz, eds., (La Jolla, California), pp. 186–189, 2003 [physics/0306116].
- [14] L. Moneta, K. Belasco, K.S. Cranmer, S. Kreiss, A. Lazzaro, D. Piparo et al., The RooStats Project, PoS ACAT (2010) 057 [1009.1003].
- [15] K. Cranmer, M. Drnevich, S. Macaluso and D. Pappadopulo, Reframing Jet Physics with New Computational Methods, EPJ Web Conf. 251 (2021) 03059 [2105.10512].
- [16] A. Güneş Baydin, L. Shao, W. Bhimji, L. Heinrich, L. Meadows, J. Liu et al., Etalumis: Bringing Probabilistic Programming to Scientific Simulators at Scale in International Conference for High Performance Computing, Networking, Storage, and Analysis (SC19), July, 2019 [1907.03382].
- [17] J. Qin and C. Tunnell, Approximate Differentiable Likelihoods for Astroparticle Physics Experiments in ACAT, 2024 [2408.09057].
- [18] A.G. Baydin, L. Shao, W. Bhimji, L. Heinrich, S. Naderiparizi, A. Munk et al., Efficient Probabilistic Inference in the Quest for Physics Beyond the Standard Model in NeurIPS, 2019, URL.
- [19] V.D. Elvira, S. Gottlieb, O. Gutsche, B. Nachman, S. Bailey, W. Bhimji et al., The Future of High Energy Physics Software and Computing in Computational Frontier Report Contribution to Snowmass 2021 [2210.05822].
- [20] K. Cranmer, G. Lewis, L. Moneta, A. Shibata and W. Verkerke, HistFactory: A tool for creating statistical models for use with RooFit and RooStats, Tech. Rep. CERN-OPEN-2012-016 (2012), DOI.
- [21] ATLAS collaboration, Likelihood preservation and statistical reproduction of searches for new physics, Tech. Rep. ATL-SOFT-PROC-2020-022 (2020), DOI.
- [22] ATLAS collaboration, Reproducing searches for new physics with the ATLAS experiment through publication of full statistical likelihoods, Tech. Rep. ATL-PHYS-PUB-2019-029 (2019).
- [23] M. Feickert, L. Heinrich and G. Stark, pyhf: a pure-Python statistical fitting library with tensors and automatic differentiation, PoS ICHEP (2022) 245 [2211.15838].
- [24] L. Heinrich, M. Feickert, G. Stark and K. Cranmer, pyhf: pure-Python implementation of HistFactory statistical models, J. Open Source Softw. 6 (2021) 2823.
- [25] L. Heinrich, M. Feickert and G. Stark, pyhf: v0.7.6, DOI.
- [26] E. Maguire, L. Heinrich and G. Watt, HEPData: a repository for high energy physics data, J. Physics: Conf. 898 (2017) 102006.
- [27] MicroBooNE collaboration, First Search for Dark-Trident Processes Using the MicroBooNE Detector, Phys. Rev. Lett. 132 (2024) 241801 [2312.13945].
- [28] Belle-II collaboration, Evidence for decays, Phys. Rev. D 109 (2024) 112006 [2311.14647].
- [29] MODE collaboration, Toward the end-to-end optimization of particle physics instruments with differentiable programming, Rev. Phys. 10 (2023) 100085 [2203.13818].
- [30] M. Feickert, L. Heinrich and M. Horstmann, Bayesian Methodologies with pyhf, EPJ Web Conf. 295 (2024) 06004 [2309.17005].
- [31] M.I. Gorinova, A.D. Gordon and C. Sutton, Probabilistic programming with densities in SlicStan: efficient, flexible, and deterministic, Proc. ACM Program. Lang. 3 (2019) 1–30 [1811.00890].
- [32] R.M. Neal, Handbook of Markov Chain Monte Carlo, Chapman and Hall/CRC (2011), DOI, [1206.1901].
- [33] M. Betancourt, A Conceptual Introduction to Hamiltonian Monte Carlo, [1701.02434].
- [34] Facebook, Prophet: Forecasting at Scale, URL.
- [35] S.J. Taylor and B. Letham, Forecasting at scale, DOI.
- [36] D. Rubin et al., Union Through UNITY: Cosmology with 2,000 SNe Using a Unified Bayesian Framework, [2311.12098].
- [37] KAGRA, VIRGO & LIGO collaboration, Population of Merging Compact Binaries Inferred Using Gravitational Waves through GWTC-3, Phys. Rev. X 13 (2023) 011048 [2111.03634].
- [38] W.M. Farr, M. Fishbach, J. Ye and D. Holz, A Future Percent-Level Measurement of the Hubble Expansion at Redshift 0.8 With Advanced LIGO, Astrophys. J. Lett. 883 (2019) L42 [1908.09084].
- [39] R. Češnovar, S. Bronder, D. Sluga, J. Demšar, T. Ciglarič, S. Talts et al., GPU-based Parallel Computation Support for Stan, [1907.01063].
- [40] J. Bradbury, R. Frostig, P. Hawkins, M.J. Johnson, C. Leary, D. Maclaurin et al., JAX: composable transformations of Python+NumPy programs, URL.
- [41] S. Maskell, Running Multiple Short MCMC Chains on a GPU Using JAX for Fast Inference with Stan in StanCon 2024 Oxford, URL.
- [42] ATLAS collaboration, Search for bottom-squark pair production with the ATLAS detector in final states containing Higgs bosons, -jets and missing transverse momentum, JHEP 12 (2019) 060 [1908.03122].
- [43] Stan Development Team, CmdStan User’s Guide, URL.
- [44] S.V. Chekanov, A note on blind technique for new physics searches in particle physics, Snowmass (2021) [2112.09548].
- [45] H. Raiffa and R. Schlaifer, Applied Statistical Decision Theory, John Wiley & Sons (1961).
- [46] Stan Development Team, Language-Specific Stan Interfaces, URL.
- [47] Stan Development Team, Stan Playground, URL.
- [48] R. Kumar, C. Carroll, A. Hartikainen and O. Martin, ArviZ: a unified library for exploratory analysis of Bayesian models in Python, J. Open Source Softw. 4 (2019) 1143.
- [49] J. Gabry and T. Mahr, bayesplot: Plotting for Bayesian Models, URL.
- [50] J. Gabry, D. Simpson, A. Vehtari, M. Betancourt and A. Gelman, Visualization in Bayesian workflow, J. R. Stat. Soc. A 182 (2019) 389 [1709.01449].
- [51] J. Gabry, R. Češnovar, A. Johnson and S. Bronder, cmdstanr: R Interface to CmdStan, URL.
- [52] J. Gabry and D. Veen, shinystan: Interactive Visual and Numerical Diagnostics and Posterior Analysis for Bayesian Models, URL.
- [53] A. Vehtari, A. Gelman, D. Simpson, B. Carpenter and P.-C. Bürkner, Rank-Normalization, Folding, and Localization: An Improved for Assessing Convergence of MCMC, Bayesian Anal. 16 (2021) [1903.08008].
- [54] R.E. Kass and A.E. Raftery, Bayes factors, J. Am. Stat. Assoc. 90 (1995) 773–795.
- [55] A. Fowlie, The Bayes factor surface for searches for new physics, Eur. Phys. J. C 84 (2024) 426 [2401.11710].
- [56] J.M. Dickey and B.P. Lientz, The Weighted Likelihood Ratio, Sharp Hypotheses about Chances, the Order of a Markov Chain, Ann. Math. Stat. 41 (1970) 214–226.
- [57] G. Cowan, K. Cranmer, E. Gross and O. Vitells, Asymptotic formulae for likelihood-based tests of new physics, Eur. Phys. J. C 71 (2011) 1554 [1007.1727].
- [58] Stan Development Team, The Bootstrap and Bagging, URL.
- [59] A.L. Read, Presentation of search results: the technique, J. Phys. G: Nucl. Part. Phys. 28 (2002) 2693–2704.